# Bernoulli shifts on additive categories and algebraic $K$-theory of wreath products

### v2

@@ 1. Introduction
is an isomorphism. The right-hand side denotes the algebraic $K$-theory groups of the group ring $R[G]$ and is of great interest for many problems in algebraic topology. The conjecture implies several other famous conjectures such as the Bass conjectures and Kaplansky’s idempotent conjecture. It has been verified for a large class of groups including all hyperbolic groups [BLR08a], CAT(0)-groups [BL12], and mapping class groups [BB19]. All these groups satisfy a more general $K$-theoretic Farrell–Jones conjecture with coefficients [BR07, Conjecture 3.2], which says that for any additive category $\mathcal{A}$ with $G$-action, the assembly map
@@
is an isomorphism. The right-hand side denotes the algebraic $K$-theory groups $K_{*}(\mathcal{A}\rtimes G)$ (see Section 2 for the definition of $\mathcal{A}\rtimes G$) where $K_{i}(-)=\pi_{i}(\mathbb{K}^{\infty}(-))$ $(i\in\mathbb{Z})$ for the non-connective algebraic $K$-theory functor $\mathbb{K}^{\infty}$ for additive categories [LS14, Definitino 4.1]. This paper is devoted to computing the algebraic $K$-theory of certain (twisted) group rings, using the Farrell–Jones conjecture as an assumption. We refer to [BLR08b] for a survey on the Farrell–Jones conjecture and to [BCH94, GAJV19] for the closely related Baum–Connes conjecture about topological $K$-theory of group $C^{*}$-algebras.
@@
Virtual equivalences induce isomorphisms at the level of $K$-theory: in the notation above, the inverse of $K_{*}(F)$ is given by $K_{*}(Q_{+})-K_{*}(Q_{-})$. Our definition allows us to prove an analogue of Izumi’s theorem [Izu19, Theorem 2.1] which says that the Bernoulli shifts of a finite group $G$ on two $KK$-equivalent $C^{*}$-algebras are $KK^{G}$-equivalent in the sense of [Kas88]. Our substitute for the $C^{*}$-algebraic $KK$-equivalences are precisely the virtual equivalences:
@@
Let $G$ be a group satisfying the $K$-theoretic Farrell–Jones conjecture with coefficients and let $H$ be a finite group. Let $R$ be a regular commutative ring such that the orders of $H$ and of every finite subgroup of $G$ are invertible in $R$. Denote by $\mathrm{FIN}(G)$ the set of finite subsets of $G$ and by $I_{R}[H]$ the augmentation ideal of $R[H]$, which is an $R$-algebra with unit by assumtion. Then we have an isomorphism
@@
We refer to Corollary 4.8 for the notation appearing in the above formula. As another application of the Going-Down principle, we compute the algebraic $K$-theory of Bernoulli shifts on many semi-simple algebras. We do this by carefully analising the $A(G)$-module structure of $K_{*}(\mathcal{A}\rtimes G)$ where $A(G)$ denotes the Burnside ring. These results are inspired by the $C^{*}$-algebraic results in [KN22, CEKN22].
@@ Acknowledgements
We would like to thank Arthur Bartels for many helpful discussions and suggestions.
@@ 2. $R$-linear categories
Throughout this paper, let $R$ be a commutative ring with unit. All tensor products appearing in this paper will be taken over $R$ and will be denoted by $M\otimes N$ rather than $M\otimes_{R}N$. An $R$-linear category is a category which is enriched over $R$-modules. An $R$-additive category is an $R$-linear category which is moreover additive. An $R$-linear functor between $R$-linear categories is a functor which is $R$-linear on morphisms sets. Note that $R$-linear functors between $R$-additive categories preserve direct sums. A basic example of an $R$-linear category is the ring $R$ itself considered as a category with one object. A basic example of an $R$-additive category is the category $\mathcal{F}^{f}(R)$ of finitely generated free $R$-modules.
@@
We introduce an equivaraint version of preferred direct sums (c.f. [LS16, Section 1.3]) which is a functorial way of adjoining directs sums. Let $\mathcal{A}$ be an $R$-linear category with $G$-action $\alpha$. We fix an infinite cardinal $\kappa$ and a $G$-set $\mathcal{U}$ such that $\mathcal{U}\times\mathcal{U}\cong\mathcal{U}$ and such that $\mathcal{U}$ contains isomorphic copies of all $G$-sets of cardinality at most $\kappa$. For instance, we can take $\kappa=|\mathbb{N}\times\mathcal{P}(G)|$ where $\mathcal{P}(G)$ denotes the power set of $G$ and
@@
where $H$ runs over all subgroups of $G$. Choose an invariant base point $u\in\mathcal{U}$ and a $G$-equivariant bijection $\tau\colon\mathcal{U}\times\mathcal{U}\to\mathcal{U}$ satisfying $\tau(u,u)=u$. We denote by $\mathcal{A}^{\omega}$ the category whose objects are tupels $(A_{z})_{z\in Z}$, where $Z\subseteq\mathcal{U}$ is a $G$-set and $A_{z}\in\mathcal{A}$ are objects. Morphisms $(A_{z})_{z\in Z}\to(B_{y})_{y\in Y}$ are given by column-finite matrices $(f_{z,y}\colon A_{z}\to B_{y})_{z\in Z,y\in Y}$ and composition is given by matrix multiplication. We equip $\mathcal{A}^{\omega}$ with the induced $G$-action $\tilde{\alpha}$ given by $\tilde{\alpha}_{g}(A_{z})_{z\in Z}\coloneqq(\alpha_{g}(A_{g^{-1}z}))_{z\in Z}$ for $(A_{z})_{z\in Z}\in\mathcal{A}^{\omega}$ and $g\in G$. For a collection $\{Z_{i}\}_{i\in I}$ of $G$-sets indexed over a $G$-set $I$, we define the graph $\mathcal{G}(\{Z_{i}\}_{i\in I})\coloneqq\{(z,i)\in\mathcal{U}\times\mathcal{U}\mid z\in Z_{i}\}$. Then $(A_{z^{\prime}})_{z=\tau(z^{\prime},i)\in\tau(\mathcal{G}(\{Z_{i}\}_{i\in I}))}$ is a concrete model for the direct sum $\bigoplus_{i\in I}(A_{z})_{z\in Z_{i}}$. We also use the suggestive notation $\bigoplus_{z\in Z}A_{z}\coloneqq(A_{z})_{z\in Z}$. We denote by $\mathcal{A}^{f}\subseteq\mathcal{A}^{\omega}$ the full subcategory of all objects $(A_{z})_{z\in Z}$ where $A_{z}\simeq 0$ for all but finitely many $z\in Z$. Note that the inclusion $\mathcal{A}\to\mathcal{A}^{f}$ induced by the choice of base-point $u\in\mathcal{U}$ is an equivalence of categories if and only if $\mathcal{A}$ is additive. We call $\mathcal{A}^{f}$ the additive completion of $\mathcal{A}$. Note that $\mathcal{A}\mapsto\mathcal{A}^{f}$ is left adjoint to the forgetful functor from the category of $R$-additive categories with $R$-linear functors to the category of $R$-linear categories together with $R$-linear functors.
@@
for $s,s^{\prime}\in S$ and $g,g^{\prime}\in G$.
@@
By a pointed category, we mean a category $\mathcal{A}$ with a preferred object $\mathds{1}_{\mathcal{A}}\in\mathcal{A}$. A functor $F\colon\mathcal{A}\to\mathcal{B}$ between pointed categories is called pointed, if it satisfies $F(\mathds{1}_{\mathcal{A}})=\mathds{1}_{\mathcal{B}}$.
@@
If $\mathcal{A}$ is a pointed $R$-linear category and if $Z$ is a (possibly infinite) set, we define the tensor product $\mathcal{A}^{\otimes Z}$ as the filtered colimit
@@
The colimit is taken inside the category of $R$-linear categories with $R$-linear functors as morphisms. Note that infinite tensor products of pointed $R$-linear categories are functorial with respect to pointed $R$-linear functors.
@@ 3. Izumi’s theorem
This section is devoted to an algebraic $K$-theory analogue of Izumi’s [Izu19, Theorem 2.1] which says that if $A$ and $B$ are $KK$-equivalent separable $C^{*}$-algebras and $G$ is a finite group, then the Bernoulli shifts $A^{\otimes G}$ and $B^{\otimes G}$ are $KK^{G}$-equivalent. Our substitute for $KK$-equivalences between $C^{*}$-algebras is the notion of virtual equivalences between additive categories introduced in Definition 2.9. For a group $G$, a $G$-set $Z$, and a (pointed) $R$-linear category $\mathcal{A}$, we always equip the tensor product $\mathcal{A}^{\otimes Z}$ with the $G$-action given by shifting the tensor factors.
@@
For any subset $T\subseteq Z$ and for any subgroup $H\subseteq G_{T}$ of the setwise stabilizer group of $T$, we define functors
@@
We construct equivariant functors $P_{\pm}\colon\mathcal{B}^{\otimes Z}\to(\mathcal{A}^{\otimes Z})^{f}$ and an equivariant natural equivalence $P_{+}F^{\otimes Z}\simeq\iota\oplus P_{-}F^{\otimes Z}$ where $\iota\colon\mathcal{A}^{\otimes Z}\to(\mathcal{A}^{\otimes Z})^{f}$ is the natural inclusion. An analogous argument will produce a natural equivalence $F^{\otimes Z}P_{+}\simeq\iota\oplus F^{\otimes Z}P_{-}$. The theorem then follows by applying $-\rtimes G$ to these equivalences and composing everything with the equivalence of $((\mathcal{A}^{\otimes Z})^{f}\rtimes G)^{f}\simeq(\mathcal{A}^{\otimes Z}\rtimes G)^{f}$ and $((\mathcal{B}^{\otimes Z})^{f}\rtimes G)^{f}\simeq(\mathcal{B}^{\otimes Z}\rtimes G)^{f}$ from Lemma 2.8.
@@
We continue this process inductively and construct $R^{(k)},n_{k},H_{i}^{(k)}$ and $S_{i}^{(k)}$ as above for $k=1,2,\dotsc,|Z|$. Now note that that $R_{H,Z}=Q_{H,Z}F^{\otimes Z}$ for any subgroup $H\subseteq G$ and that $R_{G,\emptyset}=\iota$. If we conveniently write $Q_{G,\emptyset}=Q_{H_{1}^{(0)},S_{1}^{(0)}}$ and apply the above produre until $k=|Z|$, we therefore get an equivariant natural equivalence
@@ 4. $K$-theory of wreath products
Recall that a ring $S$ is called Noetherian if every submodule of a finitely generated $S$-module is finitely generated, regular coherent if every finitely presented $S$-module has a finite projective resolution, and regular if it is Noetheriain and regular coherent [^3] In this case, every finitely generated $S$-module is automatically finitely presented.. Analogous definitions were introduced in [BL22, Definition 6.2] for any additive category $\mathcal{A}$. We recall these for convenience of the reader. A right $\mathbb{Z}\mathcal{A}$-module (or simply $\mathbb{Z}\mathcal{A}$-module) is a contravariant $\mathbb{Z}$-linear functor $\mathcal{A}\to\mathrm{Ab}$. We denote the category of $\mathbb{Z}\mathcal{A}$-modules with $\mathbb{Z}$-linear transformations as morphisms by $\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{A}$. A $\mathbb{Z}\mathcal{A}$-module $M$ is called finitely generated free if it is isomorphic to a direct sum $\bigoplus_{i\in I}\mathrm{Hom}_{\mathcal{A}}(-,A_{i})$, and finitely generated if it is isomorphic to the quotient of a finitely generated free $\mathbb{Z}\mathcal{A}$-module. A $\mathbb{Z}\mathcal{A}$-module $A$ is called finitely presented if there is an exact sequence $F_{1}\to F_{0}\to M\to 0$ where $F_{0}$ and $F_{1}$ are finitely generated free $\mathbb{Z}\mathcal{A}$-modules. A $\mathbb{Z}\mathcal{A}$-module is called finitely generated projective if it is a retract of a finitely generated free $\mathbb{Z}\mathcal{A}$-module. A finite projective resolution of a $\mathbb{Z}\mathcal{A}$-module $M$ is an exact sequence $0\to P_{n}\to\dotsc\to P_{0}\to M\to 0$ where each $P_{i}$ is a finitely generated projective $\mathbb{Z}\mathcal{A}$-module. We call $\mathcal{A}$ Noetherian if every submodule of a finitely generated $\mathbb{Z}\mathcal{A}$-module is finitely generated, regular coherent if every finitely presented $\mathbb{Z}\mathcal{A}$-module has a finite projective resolution, and regular if it is Noetheriain and regular coherent. A ring is Noetherian (resp. regular, resp. regular coherent) if and only if the category of finitely generated free (or equivalently projective) modules is [BL22, Corollary 6.5].
The following lemma is well-known for rings. We adapt the proof of [Gru08, Lemma 5.7] to the setting of additive categories. Recall that a functor between abelian categories is called flat if it preserves exact sequences.
@@
Every $\mathbb{Z}(\mathcal{A}\rtimes G)$-module $M$ is a split summand of $\iota_{*}\iota^{*}M$ in $\mathrm{Mod}\text{-}\mathbb{Z}(\mathcal{A}\rtimes G)$.
@@
(4.1) $\iota^{*}\mathrm{Hom}_{\mathcal{A}\rtimes G}(-,A)\cong\bigoplus_{g\in G}\mathrm{Hom}_{\mathcal{A}}(g(-),A)\cong\bigoplus_{g\in G}\mathrm{Hom}_{\mathcal{A}}(-,g^{-1}(A)).$

### v4

@@ 1. Introduction
is an isomorphism. The right-hand side denotes the algebraic $K$-theory groups of the group ring $R[G]$ and is of great interest for many problems in algebraic topology. The conjecture implies several other famous conjectures such as the Bass conjectures and Kaplansky’s idempotent conjecture. It has been verified for a large class of groups including all hyperbolic groups [BLR08], CAT(0)-groups [BL12], and mapping class groups [BB19]. All these groups satisfy a more general $K$-theoretic Farrell–Jones conjecture with coefficients [BR07, Conjecture 3.2], which says that for any additive category $\mathcal{A}$ with $G$-action, the assembly map
@@
is an isomorphism. The right-hand side denotes the algebraic $K$-theory groups $K_{*}(\mathcal{A}\rtimes G)$ (see Section 2 for the definition of $\mathcal{A}\rtimes G$) where $K_{i}(-)=\pi_{i}(\mathbb{K}^{\infty}(-))$ $(i\in\mathbb{Z})$ for the non-connective algebraic $K$-theory functor $\mathbb{K}^{\infty}$ for additive categories [LS14, Definition 4.1]. This paper is devoted to computing the algebraic $K$-theory of certain (twisted) group rings, using the Farrell–Jones conjecture as an assumption. We refer to [Lüc24, RV18] for surveys on the Farrell–Jones conjecture and to [BCH94, GAJV19] for the closely related Baum–Connes conjecture about topological $K$-theory of group $C^{*}$-algebras.
@@
Virtual equivalences induce isomorphisms at the level of $K$-theory: in the notation above, the inverse of $K_{*}(F)$ is given by $K_{*}(Q_{+})-K_{*}(Q_{-})$. Note that our definition can be viewed as an additive analogue of universal $K$-equivalences from [BGMN22, Definition 4.4] [^2] A subtle difference is that we don’t allow for stabilization by additional functors $S,T$ such that $FQ_{+}\oplus S\simeq\operatorname{id}_{\mathcal{B}}\oplus FQ_{-}\oplus S$ and $Q_{+}F\oplus T\simeq\operatorname{id}_{\mathcal{A}}\oplus Q_{-}F\oplus T$ as in [BGMN22, Definition 4.4]. We decided on this simplification in favor of shorter proofs rather than conceptual reasons..
Our definition allows us to prove an analogue of Izumi’s theorem [Izu19, Theorem 2.1] which says that the Bernoulli shifts of a finite group $G$ on two $KK$-equivalent $C^{*}$-algebras are $KK^{G}$-equivalent in the sense of [Kas88]. Our substitute for the $C^{*}$-algebraic $KK$-equivalences are precisely the virtual equivalences:
@@
Let $G$ be a group satisfying the $K$-theoretic Farrell–Jones conjecture with coefficients and let $H$ be a finite group. Let $R$ be a regular commutative ring such that the orders of $H$ and of every finite subgroup of $G$ are invertible in $R$. Denote by $\mathrm{FIN}(G)$ the set of finite subsets of $G$ and by $I_{R}[H]$ the augmentation ideal of $R[H]$, which is an $R$-algebra with unit by assumption. Then we have an isomorphism
@@
As another application of the Going-Down principle, we compute the algebraic $K$-theory of Bernoulli shifts on many semi-simple algebras. We do this by carefully analysing the $A(G)$-module structure of $K_{*}(\mathcal{A}\rtimes G)$ where $A(G)$ denotes the Burnside ring. These results are inspired by the $C^{*}$-algebraic results in [KN22, CEKN22].
@@ Acknowledgements
We would like to thank Arthur Bartels for many helpful discussions and suggestions. We are moreover greatful to the anonymous referee for a very detailed and helpful report.
@@ 2. $R$-linear categories
Throughout this paper, let $R$ be a commutative ring with unit. All tensor products appearing in this paper will be taken over $R$ and will be denoted by $M\otimes N$ rather than $M\otimes_{R}N$. An $R$-linear category is a category which is enriched over $R$-modules. An $R$-additive category is an $R$-linear category which is moreover additive. An $R$-linear functor between $R$-linear categories is a functor which is $R$-linear on morphism sets. Note that $R$-linear functors between $R$-additive categories preserve direct sums. A basic example of an $R$-linear category is the ring $R$ itself considered as a category with one object. A basic example of an $R$-additive category is the category $\mathcal{F}^{f}(R)$ of finitely generated free $R$-modules.
@@
We introduce an equivariant version of preferred direct sums (c.f. [LS16, Section 1.3]) which is a functorial way of adjoining direct sums. Let $\mathcal{A}$ be an $R$-linear category with $G$-action $\alpha$. We fix an infinite cardinal $\kappa$ and a $G$-set $\mathcal{U}$ such that $\mathcal{U}\times\mathcal{U}\cong\mathcal{U}$ and such that $\mathcal{U}$ contains isomorphic copies of all $G$-sets of cardinality at most $\kappa$. For instance, we can take $\kappa=|\mathbb{N}\times\mathcal{P}(G)|$ where $\mathcal{P}(G)$ denotes the power set of $G$ and
@@
where $H$ runs over all subgroups of $G$. Choose an invariant base point $u\in\mathcal{U}$ and a $G$-equivariant bijection $\tau\colon\mathcal{U}\times\mathcal{U}\to\mathcal{U}$ satisfying $\tau(u,u)=u$. We denote by $\mathcal{A}^{\kappa}$ the category whose objects are tuples $(A_{z})_{z\in Z}$, where $Z\subseteq\mathcal{U}$ is a $G$-set and $A_{z}\in\mathcal{A}$ are objects. Morphisms $(A_{z})_{z\in Z}\to(B_{y})_{y\in Y}$ are given by column-finite matrices $(f_{z,y}\colon A_{z}\to B_{y})_{z\in Z,y\in Y}$ and composition is given by matrix multiplication. We equip $\mathcal{A}^{\kappa}$ with the induced $G$-action $\tilde{\alpha}$ given by $\tilde{\alpha}_{g}(A_{z})_{z\in Z}\coloneqq(\alpha_{g}(A_{g^{-1}z}))_{z\in Z}$ for $(A_{z})_{z\in Z}\in\mathcal{A}^{\kappa}$ and $g\in G$. For a collection $\{Z_{i}\}_{i\in I}$ of $G$-sets indexed over a $G$-set $I$, we define the graph $\mathcal{G}(\{Z_{i}\}_{i\in I})\coloneqq\{(z,i)\in\mathcal{U}\times\mathcal{U}\mid z\in Z_{i}\}$. Then $(A_{z^{\prime}})_{z=\tau(z^{\prime},i)\in\tau(\mathcal{G}(\{Z_{i}\}_{i\in I}))}$ is a concrete model for the direct sum $\bigoplus_{i\in I}(A_{z})_{z\in Z_{i}}$. We also use the suggestive notation $\bigoplus_{z\in Z}A_{z}\coloneqq(A_{z})_{z\in Z}$. We denote by $\mathcal{A}^{f}\subseteq\mathcal{A}^{\kappa}$ the full subcategory of all objects $(A_{z})_{z\in Z}$ where $A_{z}\simeq 0$ for all but finitely many $z\in Z$. Note that the inclusion $\mathcal{A}\to\mathcal{A}^{f}$ induced by the choice of base-point $u\in\mathcal{U}$ is an equivalence of categories if and only if $\mathcal{A}$ is additive.
We call $\mathcal{A}^{f}$ the additive completion of $\mathcal{A}$. Our primary motivation for introducing $\mathcal{A}^{f}$ is its use in Section 3. There, given any $G$-set $Z$ and a $G$-equivariant family of functors $(F_{z})_{z\in Z}$ from $\mathcal{A}$ to $\mathcal{B}$, we will need a naturally-defined $G$-equivariant functor $\bigoplus_{z\in Z}F_{z}$ from $\mathcal{A}$ to $\mathcal{B}^{f}$.
@@
for $s,s^{\prime}\in S$ and $g,g^{\prime}\in G$.  [^4] Note that $S\rtimes G$ is precisely the result of applying the functor $\rtimes G$ to the $R$-linear category with one object and $S$ as endomorphisms.
@@
By a based category, we mean a category $\mathcal{A}$ with a preferred object $\mathds{1}_{\mathcal{A}}\in\mathcal{A}$. A functor $F\colon\mathcal{A}\to\mathcal{B}$ between based categories is called based, if it satisfies $F(\mathds{1}_{\mathcal{A}})=\mathds{1}_{\mathcal{B}}$. The strict identity condition is not essential and could be relaxed to isomorphism with minor adjustments to the subsequent discussion. However, we chose to retain the strict condition for convenience.
@@
If $\mathcal{A}$ is a based $R$-linear category and if $Z$ is a (possibly infinite) set, we define the tensor product $\mathcal{A}^{\otimes Z}$, or more precisely ($\mathcal{A},\mathds{1}_{\mathcal{A}})^{\otimes Z}$, as the filtered colimit
@@
The colimit is taken inside the category of $R$-linear categories with $R$-linear functors as morphisms. Note that infinite tensor products of based $R$-linear categories are functorial with respect to based $R$-linear functors.
@@ 3. Izumi’s theorem
This section is devoted to an algebraic $K$-theory analogue of Izumi’s [Izu19, Theorem 2.1] which says that if $A$ and $B$ are $KK$-equivalent separable $C^{*}$-algebras and $G$ is a finite group, then the Bernoulli shifts $A^{\otimes G}$ and $B^{\otimes G}$ are $KK^{G}$-equivalent. Our substitute for $KK$-equivalences between $C^{*}$-algebras is the notion of virtual equivalences between additive categories introduced in Definition 2.9. For a group $G$, a $G$-set $Z$, and a (based) $R$-linear category $\mathcal{A}$, we always equip the tensor product $\mathcal{A}^{\otimes Z}$ with the $G$-action given by shifting the tensor factors.
@@
For any subset $T\subseteq Z$ and for any subgroup $H\subseteq G_{T}$ of the setwise stabilizer group of $T$, we define $G$-equivariant functors
@@
These formulas may initially appear overwhelming, so we explain their origins. Considering the informal expression $Q_{+}-Q_{-}$ as the quasi-inverse of $F$, and leveraging the binomial formula, an initial guess for the quasi-inverse of $F^{\otimes Z}$ could be constructed by expanding $(Q_{+}-Q_{-})^{\otimes Z}$. Specifically, using a combinatorial perspective, we might propose:
$(Q_{+}-Q_{-})^{\otimes Z}=\left(\bigoplus_{\begin{subarray}{c}S\subset Z\\ \text{even}\end{subarray}}Q_{+}^{\otimes(Z\setminus S)}\otimes Q_{-}^{\otimes S}\right)-\left(\bigoplus_{\begin{subarray}{c}S\subset Z\\ \text{odd}\end{subarray}}Q_{+}^{\otimes(Z\setminus S)}\otimes Q_{-}^{\otimes S}\right),$
where the sums are taken over subsets $S\subset Z$ with even and odd cardinalities, respectively.
However, this expression turns out not to be a valid ($G$-equivariant) quasi-inverse of $F^{\otimes Z}$. The simplest counterexample arises when $G$ is the cyclic group $C_{2}$ of order 2, $Z=G$, $F=\mathrm{id}$, $Q_{+}=\mathrm{id}\oplus\mathrm{id}$, and $Q_{-}=\mathrm{id}$ on $\mathcal{A}=(\underline{\mathbb{C}})^{f}$, for example. In this case, the expansion does not satisfy the required properties of a quasi-inverse.
The core issue lies in two subtleties: the binomial formula in this setting must account for $G$-equivariant indices, which introduces additional complications when we consider the product of such formulas; and the power functor $(-)^{\otimes Z}$ is not linear, meaning it does not distribute over sums or differences in the straightforward way that one might expect.
Thus, while the proposed expansion might appear intuitive at first glance, we need a more careful treatment to correctly handle the $G$-equivariance and the nonlinearity of the power functor. The previously defined $Q_{H,T}$ accounts for a more general summand of the binomial formula, whose suitable combinations provide the correct quasi-inverse of $F^{\otimes Z}$.
@@
We construct equivariant functors $P_{\pm}\colon\mathcal{B}^{\otimes Z}\to(\mathcal{A}^{\otimes Z})^{f}$ and an equivariant natural equivalence $P_{+}F^{\otimes Z}\simeq\iota\oplus P_{-}F^{\otimes Z}$ where $\iota\colon\mathcal{A}^{\otimes Z}\to(\mathcal{A}^{\otimes Z})^{f}$ is the natural inclusion. An analogous argument will produce a natural equivalence $F^{\otimes Z}P_{+}\simeq\iota\oplus F^{\otimes Z}P_{-}$. The theorem then follows by applying $-\rtimes G$ to these equivalences and composing everything with the equivalence between $((\mathcal{A}^{\otimes Z})^{f}\rtimes G)^{f}\simeq(\mathcal{A}^{\otimes Z}\rtimes G)^{f}$ and $((\mathcal{B}^{\otimes Z})^{f}\rtimes G)^{f}\simeq(\mathcal{B}^{\otimes Z}\rtimes G)^{f}$ from Lemma 2.8.
@@
We continue this process inductively and construct $R^{(k)},n_{k},H_{i}^{(k)}$ and $S_{i}^{(k)}$ as above for $k=1,2,\dotsc,|Z|$. Now note that $R_{H,Z}=Q_{H,Z}F^{\otimes Z}$ for any subgroup $H\subseteq G$ and that $R_{G,\emptyset}=\iota$. If we conveniently write $Q_{G,\emptyset}=Q_{H_{1}^{(0)},S_{1}^{(0)}}$ and apply the above produre until $k=|Z|$, we therefore get an equivariant natural equivalence
@@ 4. $K$-theory of wreath products
Recall that a ring $S$ is called Noetherian if every submodule of a finitely generated $S$-module is finitely generated, regular coherent if every finitely presented $S$-module has a finite projective resolution, and regular if it is Noetherian and regular coherent. Analogous definitions were introduced in [BL22, Definition 6.2] for any additive category $\mathcal{A}$. We recall these for the reader’s convenience. A right $\mathbb{Z}\mathcal{A}$-module (or simply $\mathbb{Z}\mathcal{A}$-module) is a contravariant $\mathbb{Z}$-linear functor $\mathcal{A}\to\mathrm{Ab}$. We denote the category of $\mathbb{Z}\mathcal{A}$-modules with $\mathbb{Z}$-linear transformations as morphisms by $\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{A}$. A $\mathbb{Z}\mathcal{A}$-module $M$ is called finitely generated free if it is isomorphic to a direct sum $\bigoplus_{i\in I}\mathrm{Hom}_{\mathcal{A}}(-,A_{i})$, and finitely generated if it is isomorphic to the quotient of a finitely generated free $\mathbb{Z}\mathcal{A}$-module. A $\mathbb{Z}\mathcal{A}$-module $A$ is called finitely presented if there is an exact sequence $F_{1}\to F_{0}\to M\to 0$ where $F_{0}$ and $F_{1}$ are finitely generated free $\mathbb{Z}\mathcal{A}$-modules. A $\mathbb{Z}\mathcal{A}$-module is finitely generated projective if it is a retract of a finitely generated free $\mathbb{Z}\mathcal{A}$-module [^5] We encourage the reader to verify that this characterization is indeed equivalent to the definition projectivity in the sense of abelian categories.. A finite projective resolution of a $\mathbb{Z}\mathcal{A}$-module $M$ is an exact sequence $0\to P_{n}\to\dotsc\to P_{0}\to M\to 0$ where each $P_{i}$ is a finitely generated projective $\mathbb{Z}\mathcal{A}$-module. We call $\mathcal{A}$ Noetherian if every submodule of a finitely generated $\mathbb{Z}\mathcal{A}$-module is finitely generated, regular coherent if every finitely presented $\mathbb{Z}\mathcal{A}$-module has a finite projective resolution, and regular if it is Noetherian and regular coherent. A ring is Noetherian (resp. regular, resp. regular coherent) if and only if the category of finitely generated free (or equivalently projective) modules is [BL22, Corollary 6.5].
The following lemma is well-known for rings. We adapt the proof of [Gru08, Lemma 5.7] to the setting of additive categories. Recall that a functor between abelian categories is called exact if it preserves exact sequences. We call a functor $\iota\colon\mathcal{A}\to\mathcal{B}$ of additive categories flat if the induction functor $-\iota_{\ast}\colon\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{A}\to\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{B}$ is exact.
@@
The restriction functor $\iota^{*}$ is exact. The induction functor $\iota_{*}$ is exact as well, i.e. $\iota$ is flat (see below).
@@
Every $\mathbb{Z}(\mathcal{A}\rtimes G)$-module $M$ is a summand of $\iota_{*}\iota^{*}M$ in $\mathrm{Mod}\text{-}\mathbb{Z}(\mathcal{A}\rtimes G)$.
@@
(4.1) $\iota^{*}\mathrm{Hom}_{\mathcal{A}\rtimes G}(-,A)\cong\bigoplus_{g\in G}\mathrm{Hom}_{\mathcal{A}}(\alpha_{g}(-),A)\cong\bigoplus_{g\in G}\mathrm{Hom}_{\mathcal{A}}(-,\alpha_{g^{-1}}(A)).$